# Weakly-supervised diagnosis identification from Italian discharge letters

### v1

@@ 3.3 Diagnosis clustering
Table 1: Keyword-based representation of the twenty largest first-level diagnosis clusters
@@
Cluster # Cluster keywords [ITA] Cluster keywords [ENG]
Within cluster
bronch. %
@@
2 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.14% 44,48%
3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05% 11,94%
4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70,00% 15,02%
5 alimentazione bronchiolite difficolta’ feeding difficulties bronchiolitis 50.00% 2,09%
6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7% 2,39%
@@
7 broncospasmo infezione respiratoria bronchospasm respiratory infection 6.67% 0,30%
8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17% 0,60%
9 flogosi alte vie respiratorie febbre high airways flogosis fever $0$,00%
10 broncospasmo infezione vie respiratorie bronchospasm airways infection $0$,00%
11 acuto broncospasmo acute bronchospasm $0$,00%
12 broncospasmo flogosi vie bronchospasm airways flogosis $0$,00%
13 broncospasmo polmonite bronchospasm pneumonia $0$,00%
@@
14 broncospasmo corso ongoing bronchospasm $0$,00%
15 broncopolmonite sinistra left bronchopneumonia $0$,00%
16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis $0$,00%
17 broncospasmo episodio bronchospasm episode $0$,00%
18 infezione alte vie respiratorie high airways infection $0$,00%
@@
19 flogosi infezione alte vie respiratorie flogosis high airways infection $0$,00%
20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis $0$,00%
21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm $0$,00%
22 broncospasmo otite bronchospasm otitis $0$,00%
23 broncospasmo parainfettivo bronchospasm parainfective $0$,00%
@@
24 broncospasmo virosi bronchospasm virosis $0$,00%
25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory $0$,00%
26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient $0$,00%
27 altre infezioni acute vie respiratorie other acute respiratory infection airways $0$,00%
2nd level cluster # 2nd level clusters keywords [ITA] 2nd level clusters keywords [ENG] 1st level cluster # 1st level clusters keywords [ITA] 1st level clusters keywords [ENG] Within cluster bronch. % % of total bronch. covered 1 bronchiolite lieve mild bronchiolitis 1 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.16 % 51.34 % 2 bronchiolite lieve mild bronchiolitis 2 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05 % 11.94 % 3 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70.00 % 15.02 % 4 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 5 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 50.00 % 2.09 % 5 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7 % 2.39 % 6 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo infezione vie respiratorie bronchospasm infection airways 2.9 % 0.30 % 10 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo disidratazione lieve bronchospasm mild dehydration 8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17 % 0.60 % 8 broncospasmo bronchospasm 11 acuto broncospasmo acute bronchospasm 0.00 % 0.00 % 12 broncospasmo flogosi vie bronchospasm airways flogosis 13 broncospasmo polmonite bronchospasm pneumonia 14 broncospasmo corso ongoing bronchospasm 22 broncospasmo otite bronchospasm otitis 23 broncospasmo parainfettivo bronchospasm parainfective 24 broncospasmo virosi bronchospasm virosis 17 broncospasmo episodio bronchospasm episode 9 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 0.00 % 0.00 % 10 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 0.00 % 0.00 % 11 infezione vie respiratorie airways infection 18 infezione alte vie respiratorie high airways infection 0.00 % 0.00 % 19 flogosi infezione alte vie respiratorie flogosis high airways infection 0.00 % 0.00 % 27 altre infezioni acute vie respiratorie other acute respiratory infection airways 0.00 % 0.00 % 26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient 0.00 % 0.00 % 16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis 0.00 % 0.00 % 12 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 0.00 % 0.00 % 13 alte febbre flogosi respiratorie vie high fever flogosis airways 9 flogosi alte vie respiratorie febbre high airways flogosis fever 0.00 % 0.00 % 14 broncopolmonite sinistra left bronchopneumonia 15 broncopolmonite sinistra left bronchopneumonia 0.00 % 0.00 %
3.4 Disease classification
[…]
When comparing the classification results on weak labels (W) and gold labels (G), we observe a reduction in precision, affecting both the F1-Score and the AUC. However, recall is higher on gold labels, at least for BERT-based models, whereas simpler models also showed a decrease in recall.
The BERT-based models show minimal differences between settings with and without diagnosis string, indicating that the information needed to identify the diagnosis of bronchiolitis is present in the entire letter text, and these models can effectively recognize it.
As expected, the fully supervised model demonstrated better overall performance, though differences are modest, particularly in recall, which is lower on average than our best weakly-supervised model.
In Table 5, we assess the differences in the results of our best model when selecting different subsets of clusters for the weak labels. In particular, we exclude one of the 1st level clusters we selected at a time. Excluding clusters #5 or #1 does not affect the results, which is reasonable considering their smaller size. Excluding cluster #2 determines a drop in the performance and an increase in the variance, which is expected since it accounts for almost half of the bronchiolitis cases. Removing clusters #3 or #4 has a negative effect, but limited. These results confirm the solidity of our weak labels and the robustness of the model with respect to small changes in them.
Table 6 compares metrics at the hospital level, revealing notable differences. Some hospitals, such as H-0 and H-3, are more represented in the dataset, which might explain their superior performance. However, hospitals with a similar number of records show considerable performance variability, suggesting that reporting practices in discharge letters might influence disease detection. A comparison between records from pediatric vs non-pediatric ER/Departments revealed significant differences (gold AUC $80.89\pm 6.37$ vs $71.07\pm 4.57$, see Supplementary Material E for details), but there are also hospitals with pediatric ER and low performances, such as H-4. Similar findings are reported in Supplementary Material E for the different LHUs.
Table 4: 10-fold CV results for classification, comparing with rule-based baselines, fully supervised approaches and other classification modules in our pipeline. Std. Dev. in parenthesis. G = gold labels, W = weak labels.
With diagnosis string
Model
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
Umberto-E3C 82.30 (8.94) 75.46 (6.25) 77.33 (4.06) 82.30 (4.99) 72.75 (7.50) 84.59 (6.48) 75.06 (3.72) 77.68 (5.73)
Umberto 80.67 (5.99) 73.48 (7.12) 76.91 (5.97) 82.56 (4.79) 62.29 (6.24) 75.99 (4.69) 68.46 (2.77) 69.82 (4.54)
LSTM 68.57 (6.48) 64.97 (11.96) 65.54 (5.66) 67.50 (3.99) 57.39 (11.37) 60.86 (13.67) 55.08 (8.07) 63.89 (5.28)
Rule-based full text 70.18 87.05 77.70 - 65.57 68.76 67.13 -
Rule-based diagnosis 73.88 86.38 79.64 - 74.61 67.42 70.83 -
Umberto-E3C Supervised - - - - 80.30 (7.23) 81.52 (6.70) 80.91 (2.81) 86.45 (2.82)
Without diagnosis string
Model
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
Umberto-E3C 84.34 (5.19) 73.04 (10.53) 75.15 (4.27) 84.91 (3.50) 70.15 (6.40) 79.00 (5.82) 74.28 (4.10) 77.56 (3.11)
Umberto 80.01 (6.84) 68.68 (10.55) 73.90 (5.36) 80.81 (6.24) 61.41 (6.09) 76.69 (4.22) 68.18 (3.62) 70.49 (3.89)
LSTM 64.98 (6.89) 56.21 (18.88) 58.80 (13.52) 63.00 (9.24) 50.14 (8.31) 63.97 (13.52) 53.03 (9.69) 59.02 (7.36)
Rule-based full text 65.54 79.16 71.69 - 62.33 60.05 61.10 -
Umberto-E3C Supervised - - - - 75.56 (5.08) 82.55 (9.14) 78.90 (4.54) 84.05 (4.26)
Clusters for weak labels
(1st level)
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
@@ 3.4 Disease classification
Table 6: 10-fold CV results for classification, comparing different hospitals (only hospitals with more than 15 bronchiolitis cases). G = gold labels, W = weak labels.
@@
4 Discussion
This study introduces a novel pipeline that shows how weakly-supervised learning can be effectively conducted in the clinical natural language processing domain. This is the first study to address the problem of diagnosis identification from discharge letters in the Italian language, and while we validated the method using a case study on bronchiolitis, the pipeline is designed to be applicable to any disease. Our case study shows excellent results, outperforming existing unsupervised approaches, and underscores the potential for broader application in clinical settings.
We observed minimal performance differences between the metrics computed on weak and gold labels (1-2 % of F1-Score, 5-7 % of AUC), confirming that the first step of our pipeline (SAL) effectively identifies relevant cases. Despite the inevitable noise that the weak labels introduce, this is limited enough to allow the model to perform well when evaluated against gold labels. Additionally, the variations between settings where we retained the diagnosis strings and those where we removed them are negligible. This indicates that the model is effectively learning from the entire text of the discharge letters, making it particularly advantageous in situations where automatic extraction of diagnosis strings is not possible. Our manual exploration of some weak labels further revealed that diagnosis strings do not always mention the disease responsible for hospitalization. In some cases, they only refer to symptoms, and in cases of multiple diseases, not all may be reported. These findings highlight the necessity for a classification model based on the complete text of discharge letters, as the one we developed.
In contrast, rule-based approaches exhibited poor performance, mainly due to the fact that diseases like bronchiolitis are not always explicitly mentioned by name in the text. Even when mentioned, they might refer to discarded hypotheses or contacts with affected individuals, leading to false positives. While a more sophisticated rule-based system could address some of these issues, such approaches lack the scalability and generalizability of our pipeline. Conversely, our pipeline can be easily adapted to other diseases, which makes it a versatile tool for clinical NLP.
Our fine-tuned version of the Umberto model showed the best results both in the classification and the clustering modules, demonstrating a better understanding of the Italian medical domain when compared with the general-domain version.
Compared to a fully supervised method, our weakly-supervised pipeline maintained competitive performance, with only a limited and acceptable loss ($\sim 5\%$ of F1-Score, $\sim 6\%$ of AUC). This is particularly notable given the significant time and effort required for manual annotation in fully supervised setups. The pipeline has also shown robustness to the cluster selection for the weak labels. These results underscore the potential of weakly-supervised approaches in clinical natural language processing, a domain where the scarcity of labelled data is the most relevant challenge.
While our results are promising, we must acknowledge some limitations in our study that could be addressed in future work. The absence of a gold standard in labelling is a common challenge in clinical NLP, which underscores the value of weakly-supervised methods like ours. Our gold labels for bronchiolitis were manually assigned by individual physicians, making it impossible to verify inter-annotator agreement and increasing the risk of errors.
The multi-centric nature of our dataset, though a strength in its diversity, also presents an imbalance in the number of records from different hospitals and Local Health Units. Our analysis highlighted relevant performance differences among hospitals and LHUs, warranting further investigation. Testing the pipeline on a more balanced dataset would strengthen conclusions about these performance differences. Moreover, this could facilitate the development of a hierarchical transformer-based classification model that considers hospital/LHU context in the input, potentially improving model performance.
We also observed relevant differences between pediatric vs. non-pediatric ERs/Departments. Regarding this, we need to consider that pediatric admissions were identified using specific keywords. This method could miss some pediatric cases, although such instances should be limited. Future work could use data that directly includes this information as a structured field, ensuring its accuracy and completeness.
Although our case study focused on bronchiolitis, the diagnostic clusters we identified appear capable of recognizing a broad range of diseases. Validating the pipeline on other diseases would further demonstrate its robustness and generalizability. Developing a multi-disease model, particularly for conditions that share similar symptoms, could provide broader applicability and enhance the model’s utility. Tailored rule-based models often fail in such scenarios, making a machine learning-based approach more effective.
Another important direction for future research is enhancing the interpretability of our model. Identifying the portions of the discharge letter that are most relevant to the model’s output would help clinicians understand and trust the model’s decisions. Techniques related to the area of explainable AI could be employed to enhance interpretability, ensuring the model’s reasoning process is transparent and understandable. This would also help mitigate biases that could potentially be present in the training data, further enhancing the model’s reliability.
5 Conclusions
We have developed and validated the first weakly-supervised pipeline to analyze discharge letters in the Italian language, with a case study focusing on bronchiolitis. The pipeline achieved excellent results, outperforming rule-based approaches. Our fine-tuned BERT-based model showed superior performance in both the clustering and the classification module compared to a general-domain BERT model. We also assessed significant differences between hospitals/LHUs and pediatric vs. non-pediatric ERs, which suggest further investigations.
However, while the case study showcased the method’s potential, the pipeline itself is designed to be adaptable to any disease, making it a valuable tool for researchers interested in extracting diagnoses from unstructured data and for epidemiological monitoring of diseases.
Despite a small performance gap compared to fully supervised approaches, the weakly-supervised pipeline addresses real-world constraints, particularly the scarcity of labelled data in clinical NLP. Transformer-based models showed strong performance even without gold-labelled data, underscoring the viability of weakly-supervised methods in overcoming data limitations in the medical domain.
Declaration of competing interests
The authors declare that they have no competing interests.
Data statement
Data used in this study are collected and accessed through Pedianet, an Italian network of paediatricians. Data anonymization is performed in accordance with Italian regulations. Each anonymized record is assigned a unique numerical identifier. The data is then transmitted monthly to a centralized database located in Padua, Italy, for validation purposes. Inclusion in the Pedianet database is voluntary and parents/legal guardians must provide consent for their children’s anonymized data to be used for research purposes. Ethical approval of the study and access to the database were granted by the Internal Scientific Committee of So.Se.Te. Srl, the legal owner of Pedianet.
Data used in this study are not publicly available due to their confidential nature. The code developed for their analysis is available by the authors upon request.
CRediT authorship contribution statement
Acknowledgments
The present research has been partially supported by MUR, grant Dipartimento di Eccellenza 2023-2027.
Elisa Barbieri declares that her position is funded within the project “INF-ACT - One Health Basic and translational Research Actions addressing Unmet Needs on Emerging Infectious Diseases”, ID MUR PE_00000007, under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2 – CUP C93C22005170007.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT to improve the readability and clarity of some sentences. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

@@ Appendix D Additional clustering results
Table D.1: Keyword-based representation of the twenty largest first-level diagnosis clusters obtained with Umberto-TPT
@@
Table D.2: Keyword-based representation of the twenty largest first-level diagnosis clusters from the original Umberto representation
Cluster size Cluster keywords [ITA] Cluster keywords ENG
615 trauma cranico head injury
498 nato taglio born section
405 diagnosi reazione locale vaccino diagnosis specific vaccine reaction
368 infezioni specificate specified infections
333 ansia cellulite diarrea febbre frattura influenza vomito anxiety cellulite diarrhea fever fracture flu vomiting
238 aeree flogosi vie flogosis airways
232 congiuntivite specificata painful pronation
226 flogosi vie flogosis airways
206 appendicite congiuntivite enterite rinite appendicitis conjunctivitis enteritis rhinitis
199 allergica orticaria pregressa previous allergic orticaria
195 coxalgia otalgia coxalgia otalgia
178 broncospasmo episodio bronchospasm episode
178 broncospasmo laringospasmo bronchospasm laryngospasm
170 d insetto puntura insect bite
148 acuta gastroenterite acute gastroenteritis
144 alte infezione respiratorie vie high airways infection
140 enterite gastroenterite enteritis gastroenteritis
136 faringotonsillite virale viral pharyngotonsillitis
135 insorgenza recente recent onset
133 acuta bronchiolite bronchite acute bronchiolitis bronchitis
Table D.3: Keyword-based representation of the first-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset from the original Umberto representation
Cluster # Cluster keywords [ITA] Cluster keywords [ENG] Within cluster bronch. %
@@
1 acuta bronchiolite lieve severa acute mild severe bronchiolitis 87.76 12.84
2 acuta bronchiolite bronchite acute bronchiolitis bronchitis 72.93 28.96
3 bronchiolite respiratorio virus bronchiolitis respiratory virus 66.66 2.99
4 bronchiolite positiva vrs bronchiolitis rsv positive 44.44 1.19
5 bronchiolite complicata corso ex gastroenterite paziente bronchiolitis complicated course former gastroenteritis patient 40.00 0.60
6 bronchite broncospasmo lieve bronchitis mild bronchospasm 20.00 0.30
7 bronchiolite broncospasmo lieve bronchiolitis mild bronchospasm 16.67 0.30
8 allarme bronchiolite caratteristiche d epistassi insetto locale lussazione medicazione pregressa reazione sdr sn vrs bronchiolitis alarm features epistaxis local insect dislocation dressing previous reaction sdr sn rsv 16.67 0.30
9 acuta insufficienza respiratoria acute respiratory failure 14.04 2.39
10 broncospasmo respiratoria bronchospasm respiratory 12.50 2.99
11 broncospasmo laringospasmo bronchospasm laryngospasm 1.12 0.60
12 broncospasmo episodio bronchospasm episode 0.56 0.30
13 febbrile respiratoria respiratory fever 0.0 0.0
14 febbre recente respiratorie vie recent fever airways 0.0 0.0
15 corso infezione respiratoria ongoing respiratory infection 0.0 0.0
16 broncospasmo vie bronchospasm airways 0.0 0.0
17 broncospasmo paziente ricorrente bronchospasm recurrent patient 0.0 0.0
18 broncospasmo parainfettivo parainfective bronchospasm 0.0 0.0
19 broncospasmo noto known bronchospasm 0.0 0.0
20 broncospasmo infettivo parainfettivo parainfective infectious bronchospasm 0.0 0.0
21 broncospasmo corso virosi bronchospasm ongoing virosis 0.0 0.0
22 broncospasmo corso respiratorie vie bronchospasm ongoing airways 0.0 0.0
@@
23 broncospasmo corso infezione respiratoria bronchospasm ongoing respiratory infection 0.0 0.0
24 broncospasmo corso flogosi vie bronchospasm ongoing flogosis airways 0.0 0.0
@@
25 broncospasmo corso flogosi lieve respiratoria bronchospasm ongoing mild flososis respiratory 0.0 0.0
26 broncopolmonite destra iniziale initial right bronchopneumonia 0.0 0.0
27 batterica broncopolmonite esterna infettiva otite polmonite specificata bacterial bronchopneumonia external infectious otitis pneumonia specified 0.0 0.0
28 alte vie respiratorie infezione high airways infection 0.0 0.0
29 alte vie respiratorie flogosi high airways flogosis 0.0 0.0
@@
30 corso alte vie respiratorie flogosi ongoing high airways flogosis 0.0 0.0
31 corso alte vie respiratorie virosi febbre ongoing high airways virosis fever 0.0 0.0
32 acuto broncospasmo lieve moderato acute mild moderate bronchospasm 0.0 0.0
33 acute infezioni vie respiratorie superiori acute high airways infection 0.0 0.0
34 acuta alte vie respiratorie rinite virosi acute high airways virosis rhinitis 0.0 0.0
Table D.4: Keyword-based representation of the second level clusters related to respiratory infections from the original Umberto representation
2nd level cluster # 2nd level clusters keywords [ITA] 2nd level clusters keywords [ENG] 1st level cluster # 1st level clusters keywords [ITA] 1st level clusters keywords [ENG]
1 acuta corso infezione insufficienza respiratoria acute bronchiolitis undergoing respiratory failure # 9 acuta insufficienza respiratoria acute respiratory failure
# 15 corso infezione respiratoria ongoing respiratory infection
2 broncospasmo episodio noto vie known bronchospasm episode airways # 16 broncospasmo vie bronchospasm airways
# 19 broncospasmo noto known bronchospasm
# 12 broncospasmo episodio episode bronchospasm
@@
3 broncospasmo corso paziente ricorrente virosi bronchospasm ongoing recurrent patient virosis # 21 broncospasmo corso virosi bronchospasm ongoing virosis
# 17 broncospasmo paziente ricorrente bronchospasm recurrent patient
4 broncospasmo lieve bronchospasm mild # 6 bronchiolite broncospasmo lieve bronchiolitis bronchospasm mild
# 7 bronchiolite broncospasmo lieve bronchiolitis bronchospasm mild
# 32 acuto broncospasmo lieve moderato acute bronchospasm mild moderate
5 acuta bronchiolite acute bronchiolitis # 2 acuta bronchiolite bronchite acute bronchiolitis bronchitis
# 1 acuta bronchiolite lieve severa acute bronchiolitis mild severe
6 respiratorie vie airways # 28 alte infezione respiratorie vie high airways infection
# 31 alte vie respiratorie corso febbre virosi high airways ongoing fever virosis
# 33 acute infezioni vie respiratorie superiori acute high airways infection
# 34 acuta alte respiratorie rinite vie virosi acute high airways rhinitis virosis
# 14 febbre recente respiratorie vie recent fever airways
8 broncospasmo respiratoria bronchospasm respiratory # 23 broncospasmo corso infezione respiratoria bronchospasm ongoing respiratory infection
# 25 broncospasmo corso flogosi lieve respiratoria bronchospasm ongoing flogosis mild respiratory
# 10 broncospasmo respiratoria bronchospasm respiratory
9 broncospasmo corso vie bronchospasm ongoig airways # 24 broncospasmo corso flogosi vie bronchospasm ongoing flogosis
# 22 broncospasmo corso respiratorie vie bronchospasm ongoing airways
10 alte flogosi respiratorie vie flogosis high airways # 30 corso flogosi alte vie respiratorie flogosis high airways ongoing
# 29 flogosi alte vie respiratorie flogosis high airways
11 bronchiolite respiratorio virus bronchiolitis respiratory virus # 3 bronchiolite respiratorio virus bronchiolitis respiratory virus
12 bronchiolite positiva vrs bronchiolitis rsv positive # 4 bronchiolite positiva vrs bronchiolitis rsv positive
13 bronchiolite complicata corso ex gastroenterite paziente bronchiolitis complicated course former gastroenteritis patient # 5 bronchiolite complicata corso ex gastroenterite paziente bronchiolitis complicated course former gastroenteritis patient
14 allarme bronchiolite caratteristiche d epistassi insetto locale lussazione medicazione pregressa reazione sdn sn vrs bronchiolitis alarm features epistaxis local insect dislocation dressing previous reaction sdn sn rsv # 8 allarme bronchiolite caratteristiche d epistassi insetto locale lussazione medicazione pregressa reazione sdn sn vrs bronchiolitis alarm features epistaxis local insect dislocation dressing previous reaction sdn sn rsv
15 febbrile respiratoria respiratory fever # 13 febbrile respiratoria respiratory fever
16 broncopolmonite destra iniziale initial right bronchopneumonia # 26 broncopolmonite destra iniziale initial right bronchopneumonia
17 batterica broncopolmonite esterna infettiva otite polmonite specificata bacterial bronchopneumonia external infectious otitis pneumonia specified # 27 batterica broncopolmonite esterna infettiva otite polmonite specificata bacterial bronchopneumonia external infectious otitis pneumonia specified
Appendix E Additional classification results
In this section, we report additional classification results to assess the effect and the differences among hospitals and LHUs, in Tables E.1 and E.2, respectively. Since the number of cases for each hospital and LHU is limited, we consider those with a minimum number of bronchiolitis cases (15) and measure the results using a Leave One Group Out (LOGO) approach. For each hospital/LHU to be evaluated, we train the model on the data from the other hospitals/LHUs, and we measure the performance on the data from the hospital/LHU of interest.
In Table E.3 we also compare results between the groups of pediatric vs non-pediatric ER/Department admissions. In this case, we adopt the same 10-fold stratified cross-validation used to compare different classification models, dividing the test set between pediatric and non-pediatric ERs/Departments at each iteration and computing the metrics on these two groups separately.
We identified the records related to pediatric ERs/Departments by looking at the presence of specific keywords (pediatria, pediatrico, pediatrica) in the header of the letters. This marked 10’340 records (31.17%) as related to pediatric ERs/Departments.
@@ Appendix E Additional classification results
Table E.1: 10-fold CV results for classification, comparing different hospitals (only hospitals with more than 15 bronchiolitis cases). G = gold labels, W = weak labels.
@@
Table E.2: 10-fold CV results for classification, comparing different LHUs (only LHUs with more than 15 cases). Std. Dev. in parenthesis. G = gold labels, W = weak labels.
LHU
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
LHU-3 75.56 65.38 70.10 78.81 69.70 53.49 60.53 73.94
LHU-5 83.33 40.00 54.05 65.77 45.83 64.71 53.66 58.25
LHU-2 94.12 64.00 76.19 88.89 51.61 94.12 66.67 59.95
LHU-6 64.29 81.81 72.00 91.17 51.67 71.43 59.96 83.52
LHU-7 86.27 83.02 84.62 87.06 82.50 76.74 79.52 86.22
LHU-0 77.46 78.57 78.01 90.40 81.40 63.64 71.43 86.56
Table E.3: 10-fold CV results for classification, comparing records from pediatric ERs/Departments with records from non-pediatric ERs/Departments. Std. Dev. in parenthesis. G = gold labels, W = weak labels.
Pediatrics ER
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
Yes 87.79 (7.12) 83.31 (4.30) 85.13 (6.56) 86.04 (5.29) 75.89 (9.85) 88.57 (8.23) 81.74 (7.71) 80.89 (6.37)
No 72.47 (7.28) 80.52 (7.87) 76.08 (5.64) 75.93 (4.04) 60.29 (8.29) 85.21 (6.11) 69.37 (4.36) 71.07 (4.57)